### v1

# you are in the wrong room

Abstract.
If taken seriously, the advice in the title leads to interesting combinatorics. Consider $N$ people moving between $M$ rooms as follows: at each step, simultaneously, the smartest person in each room moves to a different room of their choice. There are $M^{N}$ possible configurations. We classify the pairs $(N,M)$ such that every configuration can be reached from every other. More precisely, let $G(N,M)$ be the directed graph with vertices representing configurations and edges representing legal moves. We prove that the graph $G(N,M)$ is strongly connected if and only if $M\geq N+1$ (one extra room for maneuvering is both required and sufficient). In the case $M\leq N$, we explicitly describe the components. Proofs are based on algebraic properties of derangements.
1991 Mathematics Subject Classification:
Primary 05C20, 05C57
1. Introduction
Fix $N\geq 1$ and $M\geq 2$. Denote $[K]=\{1,\ldots,K\}$. Let $V=\{f:[N]\to[M]\}$ and let $E$ be the set of pairs $(f,g)\in V\times V$ with the following property: for all $k\in[N]$ we have $g(k)\neq f(k)$ if and only if $k=\max f^{-1}(f(k))$. Define $G(N,M)=(V,E)$.
Interpreted as in the abstract, the total order $1<\cdots<N$ represents smartness. The functions $f$ and $g$ represent two possible configurations of $N$ people in $M$ rooms. The property “$g(k)\neq f(k)$ if and only if $k=\max f^{-1}(f(k))$” states that, from the configuration $f$ to $g$, those changing room are precisely the smartest individuals in their respective room.
It is natural to ask whether any configuration can be reached from any other in a finite number of steps. The following theorem shows that having one extra room for maneuvering is both necessary and sufficient.
Theorem 1.1.
The graph $G(N,M)$ is strongly connected if and only if $M\geq N+1$.
By Theorem 1.1, for $M\leq N$ the graph $G(N,M)$ is not strongly connected. The next theorem shows that, in such case, a positive fraction of vertices forms a single strongly connected component (a giant component), while the remaining vertices have zero in-degree and positive out-degree. In other words, every directed path enters the giant component in at most one step and never leaves it afterwords.
Theorem 1.2.
Suppose that $3\leq M\leq N$. Then $G(N,M)$ is weakly connected, has $1$ strongly connected component of size $M^{N}-M^{N-M+1}$, and $M^{N-M+1}$ strongly connected components of size $1$, each containing a zero in-degree vertex, representing the configurations in which the $M$ smartest people share the same room.
Figure 1. The configuration graph $G(4,3)$ for $N=4$ people and $M=3$ rooms. The set $V_{s}$ is labeled $\bullet$, the set $V_{c}$ is labeled $\blacksquare$, the other vertices are labeled $\times$. The vertices labeled $\bullet$ and $\times$ form a single strongly connected component. Every vertex labeled $\blacksquare$ has zero in-degree and form a strongly connected component on its own.
The case $2=M<N$, excluded from Theorem 1.2, is exceptional but is easy to describe: the two smartest people occupy both rooms in at most one step, and swap places with no-one else moving afterwards. In this case $G(N,M)$ is union of small weakly connected components.
The proofs of Theorem 1.1 and Theorem 1.2 are explicit and construct paths of length $\mathcal{O}(N)$. In general, this cannot be improved: if all people start in the same room, it takes at least $N$ steps for all of them to change room.
We conclude the introduction with an open problem. Suppose each room’s smartest person independently choses the next room uniformly at random among the other rooms. Then the next vertex in $G(N,M)$ is chosen uniformly at random among the out-neighbors. While Theorem 1.1 and Theorem 1.2 classify recurrent vertices in $G(N,M)$, they do no determine their asymptotic frequencies. This is left as an open problem. For a survey on random walks on graphs, see [4]. Note that the random walk on $G(N,M)$ can also be viewed as $N$ random walkers moving among $M$ rooms under a priority rule. While multiple random walks have been extensively studied [1], such a priority rule seems to be novel.
2. Proof of Theorem 1.1
To prove Theorem 1.1, we need to show that for every $f,g\in V$ there is a directed path from $f$ to $g$. We will do so by describing a subset $V_{s}\subseteq V$ that is reachable from every $f\in V$ and that can reach every $g\in V$. We will also describe a subset $V_{c}\subseteq V$ with zero in-degrees. The proof of Theorem (1.1) is broken down in a series of lemmas that will be needed again for Theorem 1.2. As in the introduction, we assume $N\geq 1$ and $M\geq 2$ to avoid trivial cases (for $M=1$ the problem is ill-defined since the smartest person has nowhere to go).
Inspired by [6], we use brackets to denote integer intervals. For $K,H\in\mathbb{Z}$ let
$[K,H]=\{K,K+1,\ldots,H\}\cap\mathbb{Z}_{\geq 1}$
and $(K,H]=[K,H]\setminus\{K\}$. If $1\leq K\leq H$ then $[K,H]=\{K,K+1,\ldots,H\}$. If $K\leq 1\leq H$ then $[K,H]=[H]$. With this notation $(N-M,N]$ represents the $M$ smartest people among $1,\ldots,N$: if $M\geq N$, it includes all of them.
Definition 2.1.
Let $V_{s}\subseteq V$ be the set of configurations $f$ such that the restriction of $f$ to $(N-M,M]$ is injective.
Intuitively $V_{s}$ is the set of configurations in which the $M$ smartest people among $1,\ldots,N$ occupy different rooms; if $N\leq M$, this means that everyone is in a different rooms. As we will see, the set $V_{s}$ serves as a central hub in the graph $G(N,M)$, see Figure 1.
Lemma 2.2.
The set $V_{s}$ can be reached from any other configuration in at most $N$ steps.
Proof.
Note that the movements of $[1,N-M]$ do not affect those of $(N-M,M]$. Therefore we can ignore $[1,N-M]$ and suppose $N\leq M$.
Fix $f\in V$. Let $g$ be obtained from $f$ as follows: for each room with at least $2$ people move the smartest person to a distinct empty room, and derange the set of people who occupy a room individually. Then $(f,g)$ is an edge. Unless $f\in V_{s}$, the configuration $g$ occupies a strictly greater number of rooms. By proceeding this way, we reach $V_{s}$ in at most $N$ steps. ∎
Derangements are permutations without fixed points. The following lemma shows that, with the exception of $n=3$, the symmetric group $\mathrm{Sym}(n)$ is generated by derangements. The case $n=3$ is indeed an exception since derangements of $3$ objects are even permutations.
Lemma 2.3.
For every integer $n\neq 3$ the symmetric group $\mathrm{Sym}(n)$ is generated by derangements. More precisely, every permutation of $n\neq 3$ objects is product of at most $4$ derangements.
Proof.
The statement is vacuous for $n=1$ and trivial for $n=2$. Let $n\geq 4$. Husemoller [3, Proposition 4] attributes to Gleason the result that every even permutation is product of two $n$-cycles (see [5] for a more general result). In particular, every even permutation is product of $2$ derangements. Since every odd permutation is product of an even permutation and a transposition, it remains to show that every transposition is product of $2$ derangements.
We will use the standard cycle notation of permutations [2]. Consider the transposition $\tau=(a_{2}a_{4})$ and the $n$-cycle $\sigma=(a_{1}\cdots a_{n})$. Then
$\tau\circ\sigma=(a_{1}a_{4}\cdots a_{n})\circ(a_{2}a_{3})$
is a derangement. Clearly $\sigma^{-1}$ is also a derangement. Therefore the transposition $(a_{1}a_{3})=\sigma^{-1}\circ\tau\circ\sigma$ is product of two derangements. ∎
The next lemma shows that if $f,g\in V_{s}$ differ only on $(N-M,M]$, there is a directed path from $f$ to $g$. In the case $M\neq 3$ the path lies in $V_{s}$, while the case $M=3$ requires leaving $V_{s}$ momentarily.
Lemma 2.4.
Let $f,g\in V_{s}$. Suppose that the restrictions of $f$ and $g$ to $[1,N-M]$ are equal. Then $g$ can be reached from $f$ in at most $4$ steps.
Proof.
If $M\neq 3$ this follows from Lemma 2.3: since $g\mid_{(N-M,N]}=\sigma\circ f_{(N-M,N]}$ for some permutation $\sigma\in\mathrm{Sym}([M])$, it can be reached in at most $4$ steps.
Let $M=3$. Since derangements of $3$ objects are $3$-cycles, it is enough to prove the case in which $g$ is obtained from $f$ by a transposition. The case $N=2$ is trivial. Let $N=M=3$. Up to relabeling rooms, we can suppose $f=(f(1),f(2),f(3))=(1,2,3)$ and $g=(2,1,3)$. The following is a directed path from $f$ to $g$:
$(1,2,3),(2,3,2),(2,2,1),(2,1,3).$
Now let $3=M<N$. Up to relabeling rooms and up to rotating $(N-M,M]$, we can suppose $f=(\ldots,1,1,2,3)$ and $g=(\ldots,1,2,1,3)$, where $\ldots$ is the same for both and will not be changed. The following is a directed path from $f$ to $g$:
$(\ldots,1,1,2,3),(\ldots,1,2,3,2),(\ldots,3,2,2,1),(\ldots,1,2,1,3).$
This concludes the proof. ∎
Definition 2.5.
Let $V_{c}\subseteq V$ be the set of configurations $f$ such that the restriction of $f$ to $(N-M,M]$ is constant.
Intuitively $V_{c}$ is the set of configurations in which the $M$ smartest people among $1,\ldots,N$ occupy the same room: if $N\leq M$, this means that all people are in the same room. As we will see, while the set $V_{s}$ is a central hub in the graph $G(N,M)$, the set $V_{c}$ is poorly connected, see Figure 1.
Lemma 2.6.
Let $M\leq N$. Then there is no edge $(f,g)$ with $g\in V_{c}$.
Proof.
As in Lemma 2.2, note that the movements of $[1,N-M]$ do not affect those of $(N-M,M]$. Therefore we can ignore $[1,N-M]$ and suppose $N=M$.
Let $g(k)=r$ for all $k\in[N]$ and suppose that there is an edge $(f,g)\in E$. The set $f^{-1}(r)$ is empty: otherwise $k=\max f^{-1}(r)$ would move to a different room $g(k)\neq r$. Moreover, for every room $s\neq r$ we have $\left|f^{-1}(s)\right|=1$: otherwise $k=\min f^{-1}(s)$ would remain in the same room $g(k)=s\neq r$. Therefore $f$ injects $[N]$ into $[M]\setminus\{r\}$, contradicting $M\leq N$. ∎
The previous lemma shows that vertices in $V_{c}$ have zero in-degree. The following lemma implies that vertices in $V\setminus V_{c}$ have positive in-degree.
Lemma 2.7.
Let $g\in V\setminus V_{c}$. Then there is $f\in V_{s}$ such that $(f,g)$ is an edge.
Proof.
Let $A\subseteq(N-M,N]$ the set of people among $(N-M,N]$ that are the smartest in some room according to $g$. Since $g\notin V_{c}$, this set has at least $2$ elements, in particular there exists a derangement $\sigma$ of $A$. Let $\iota$ be any bijection from $(N-M,N]\setminus A$ to $[M]\setminus g(A)$. Let $f\in V$ be defined as follows:
[…]
Then $f\in V_{s}$ and $(f,g)$ is an edge. ∎
Proof of Theorem 1.1.
First, suppose that $M\geq N+1$. We will prove that $G(N,M)$ is strongly connected. Since $M\geq N+1$ the set $V_{s}$ is the subset of injective functions and $V_{c}$ the subset of constant functions. Fix $f,g\in V$. By Lemma 2.2, there is a directed path of length at most $N$ from $f$ to some $f_{1}\in V_{s}$. By Lemma 2.7 there is a directed path of length $1$ from some $g_{1}\in V_{s}$ to $g$. Since $[1,N-M]=\emptyset$, the restrictions of $f_{1}$ and $g_{1}$ to $[1,N-M]$ are equal (they are both empty functions), thus by Lemma2.4 there is a path of length at most $4$ from $f_{1}$ to $g_{1}$. We conclude that every $g$ can be reached from every $f$ in $\mathcal{O}(N)$ steps.
Conversely, suppose that $G(N,M)$ is strongly connected. Then there is an edge $(f,g)$ with $f\in V$ and $g\in V_{c}$. By Lemma 2.6 this implies $M\geq N+1$. ∎
3. Proof of Theorem 1.2
In order to prove Theorem 1.2, it is convenient to first consider a simplified version of the problem in which only one person moves at a time.
Lemma 3.1.
Let $N\geq 1$ and $M\geq 2$. Consider a variant of our setup in which, at each step, a room is selected and only the smartest person in that room moves to another room, while no one else moves. Let $G^{\prime}(N,M)$ be the corresponding configuration graph. The graph $G^{\prime}(N,M)$ is strongly connected. In particular, any vertex can be reached from any other in $\mathcal{O}(N)$ steps.
Proof.
The statement is trivial for $N=1$. For $N\geq 2$, begin by moving one at a time all individuals from the room occupied by the person $1$. Once $1$ is alone, move $1$ to the intended room. Once $1$ is settled, it is never forced to move and it does not affect the movement of $[2,N]$. The statement follows by induction. ∎
The idea behind the following lemma is as follows: the movements of $[1,N-M]$ allowed by $G^{\prime}(N-M,M)$ can be achieved in $G(N,M)$, although in a slightly larger number of steps, while starting and ending in $V_{s}$.
Lemma 3.2.
Let $2\leq M\leq N$. Let $h_{1},h_{2}:[1,N-M]\to[M]$. For every $f_{1}\in V_{s}$ with $f_{1}\mid_{[1,N-M]}=h_{1}$ there is $f_{2}\in V_{s}$ with $f_{2}\mid_{[1,N-M]}=h_{2}$ and a directed path in $G(N,M)$ from $f_{1}$ to $f_{2}$ of length $\mathcal{O}(N)$.
Proof.
By Lemma 3.1, we can suppose that there is a unique $j\in[1,N-M]$ such that $h_{1}(j)\neq h_{2}(j)$ and $j=\max h_{1}^{-1}(h_{1}(j))$. The general case follows by induction.
Let $r=h_{1}(j)$. Let $f_{1}\in V_{s}$ such that $f_{1}\mid_{[1,N-M]}=h_{1}$. We are going to define $g,f_{2}\in V$ such that
  1. (i)
    $(f_{1},g)$ and $(g,f_{2})$ are edges of $G(N,M)$;
  2. (ii)
    $\max g^{-1}(r)=j$;
  3. (iii)
    $f_{2}\in V_{s}$ and $f_{2}\mid_{[1,N-M]}=h_{2}$.
Note that in the configuration $f_{1}$ the person $j$, defined above, is the second smartest person in the room $r$. The smartest person in $r$ is some $k\in(N-M,N]$.
Let $\rho$ be a derangement of $[M]\setminus\{r\}$. Let $g$ be such that $s=g(k)\neq r$, such that
$g\mid_{(N-M,N]\setminus\{k\}}=\rho\circ(f_{1}\mid_{(N-M,N]\setminus\{k\}}),$
and such that $g(i)=f_{1}(i)$ for every $i\in[1,N-M]$. Note that in the configuration $g$ the person $j$ is the only person among $[1,N-M]$ that can (and has to) move in the next step. Moreover, note that $s$ is the only room containing more than $1$ (exactly $2$) people from $(N-M,M]$; in particular $n=\min g^{-1}(s)\cap(N-M,M]$ is the only person among $(N-M,N]$ that cannot (and does not) move next step.
Let $\sigma$ be a derangement of $[M]$ mapping $r$ to $s$. Let $f_{2}$ be such that $f_{2}(j)=h_{2}(j)$, such that $f_{2}(i)=g(i)=h_{2}(i)$ for every $i\in[1,N-M)$, such that $f_{2}(n)=s$, and such that
$f_{2}\mid_{(N-M,N]\setminus\{n\}}=\sigma\circ(g\mid_{(N-M,N]\setminus\{n\}}).$
The configurations $g$ and $f_{2}$ have the desired property. ∎
Proof of Theorem 1.2.
Let $3\leq M\leq N$. Lemma 2.6 shows that every $g\in V_{c}$ has zero in-degree. Every $g\in V_{c}$ forms strongly connected component of its own and has positive out-degree. Therefore, it is enough to prove that $V\setminus V_{c}$ is a strongly connected component.
Let $f,g\in V\setminus V_{c}$. We need to prove that there is a directed path from $f$ to $g$. By Lemma 2.2, there is a directed path of length at most $N$ from $f$ to some $f_{1}\in V_{s}$. By Lemma 2.7, there is a directed path of length $1$ from some $g_{1}\in V_{s}$ to $g$. In contrast to the proof of Theorem 1.1, we cannot apply Lemma 2.4 yet: since $M\leq N$ the restrictions $h_{1}=f_{1}\mid_{[1,N-M]}$ and $h_{2}=g_{1}\mid_{[1,N-M]}$ might differ. By Lemma 3.2 there is $f_{2}\in V_{s}$ such that $f_{2}\mid_{[1,N-M]}=h_{2}$ and such that $f_{2}$ can be reached from $f_{1}$ in $\mathcal{O}(N)$ steps. Lemma 2.4 can now be applied to $f_{2}$ and $g_{1}$. We conclude that every $g\in V\setminus V_{c}$ can be reached from every $f\in V\setminus V_{s}$ in $\mathcal{O}(N)$ steps. ∎